# Quantum walks and entanglement in cavity networks

### v1

@@ V.1 Entanglement witnesses
One experiment-friendly way to assess entanglement is given in terms of measurable operators, dubbed entanglement witnesses [83, 84, 85, 86]. One way to construct such witnesses is based on finding the maximal (likewise, minimal) expectation values of an observable $\hat{L}$ for separable states via generalized eigenvalue problems [87, 88], so-called separability eigenvalue equations. The thereby constructed witnesses are directly applicable to detecting multipartite entanglement in experiments [89, 90, 91].
@@
where $\hat{\rho}_{\text{sep}}$ is a non-entangled state and $\mathcal{S}$ denotes the set of all pure separable states. We have indicated with $g_{\max}$ the largest possible expected value of $\hat{L}$ for separable states. This maximal expectation value is identical to the maximal separability eigenvalue from the construction approach mentioned before [87, 88].
@@
where $\lambda_{\max}$ denotes the maximal expected value for all separable and inseparable states—i.e., the maximal ordinary eigenvalue of $\hat{L}$—and is, for all examples considered here, $\lambda_{\max}=1$. Thus, $E$ is yields a value between zero and one and is zero for separable states, hence quantifying the witnessed entanglement [92].
@@
where the phases $\varphi_{k}$ are adjusted such that they coincide with the single-excitation state under study for maximizing the expectation value $\langle\hat{L}\rangle$. Regardless the phases, the maximum expected value for separable states, i.e., maximal separability eigenvalue, reads [93, 94]
@@
With the observable described thus far, we probe how similar the system’s state is to an $N$-partite, maximally multipartite-entangled state in a $W$ configuration that includes one excitation. For the states produced in our quantum walk scenario, Eqs. (9) and (12), we can directly compute the sought-after expectation values. Since we have $N=N_{1}N_{2}$ atoms and cavities for the considered geometries, for the cavity-only entanglement, we find
@@
also yielding $\lambda_{\max}=1$. Note that $|\vec{0}\rangle$ is free of excitations and, thus, perpendicular to the single-excitation cavity state $|\vec{\alpha}\rangle$ and the single-excitation atomic state $|\vec{\beta}\rangle$. As $|\Phi\rangle$ is thus given in its Schmidt decomposition, the bipartite cavity-atom separability bound $g_{\max}$ can readily found via the maximal Schmidt coefficient [87],
@@
Figure 4: Evolution of the entanglement in the quantum walk for a network with toroidal structure (top row) and a Möbious strip configuration (bottom row) in terms of the entanglement quantifier in Eq. (22). Lines (in color) represent the ideal, lossless entanglement evolution ($\Gamma^{(C)}/\kappa=\Gamma^{(A)}/\kappa=0$) whereas decaying curves (in gray) show attenuated situations ($\Gamma^{(C)}/\kappa=\Gamma^{(A)}/\kappa=0.0025$). The multipartite entanglement of the full system is shown in panels (a1) and (b1) for the two geometries. The multipartite entanglement between the cavity modes alone is shown in (a2) and (b2), and the analogous case for atoms in (a3) and (b3). In (a4) and (b4), the bipartite light-matter entanglement is shown. The other parameters—in addition to the decay rates—are the same as in Fig. 3.
@@ V.2 Entanglement evolution
The evolution of the aforementioned four kinds of entanglement is displayed in Fig. 4 for the torus (top row) and Möbius strip (bottom row). In the four depicted scenarios, using $E$ as given in Eq. (22), we consider both a lossless evolution (color) and an evolution subject to losses (gray). The multipartite entanglement of the full system is shown in the first column. This includes all quantum correlations between the $2N$ elements (all cavity modes and atoms) and is $E=1$ for a maximal entanglement in a $W$-state configuration of the single excitation. The second column similarly depicts the entanglement between all cavity fields alone, being the $W$ entanglement of the system after tracing over the atomic degrees of freedom. Likewise, the entanglement between the atoms is shown in the third column. Finally, the bipartite light-matter entanglement, i.e., the entanglement between the collection of all atoms and the collection of all cavity fields, is presented in panels (a4) and (b4).
When there are no losses in the system, we can see in the fourth column how the entanglement between light and matter is constant and maximum, $E(t)=1$, except for the initial time, $E(t_{0}=0)=0$, as we prepare the system in the separable configuration $|\bar{\psi}(t_{0})\rangle=|1_{k}^{(C)}\rangle$, with $k=(3,3)$. This discontinuous behavior reflects the fact that as soon as the evolution of the initial photon excitation gives a non-zero probability of finding an atomic excitation, the light and matter subsystems become entangled. When losses are included, however, these perfect correlations decay, as seen in panels (a4) and (b4).
The multipartite entanglement of the full system, panels (a1) and (b1) in Fig. 4, is the most robust kind of entanglement against losses since it is the one with the highest degree of separation, i.e., each atom and cavity mode is considered as one subsystem. Note that, even if concrete values are not identical for both geometries, the general behavior is analogous for the quantum walks on torus and Möbius-strip graphs, including the time decay due to losses, for all kinds of entanglement. Furthermore, we observe several instances of increasing and decreasing entanglement, including so-called sudden deaths and sudden births [95]. In general, for the four kinds of entanglement analyzed here, one can observe that a considerable amount of entanglement is created during the evolution and sustained over time, as described by the envelopes of $E(t)$ and notwithstanding short-time oscillations within.
@@ VI Conclusions
The performance of a multipartite light-matter network with arbitrary configurations as a quantum walk has been considered. The quantum dynamics and the multipartite entanglement have been analyzed, focusing on two highly nontrivial geometries: a torus as a genus-one topology and a Möbius strip as a non-orientable manifold. Clear signatures of quantumness—distinct from the dispersive random-walk evolution—were observed in our computation, which was based on the exact solutions we derived. Also, the impact of optical losses and spontaneous, atomic decays were included in our model.
Quantum interference patterns in the probability distributions were found for both geometries. This included oscillatory features in the optical part, in the atomic part, as well as between them. A comparison of random walks and quantum walks further showed that the quantum interference is highly sensitive to the underlying geometry while no distinctive features were observable in the classical diffusion.
Additional quantum properties of the network are confirmed by the entanglement created through the open-system evolution. The analysis of entanglement includes the multipartite entanglement of the full system, the multipartite entanglement between the atoms, the multipartite entanglement between the cavity modes, and the bipartite light-matter hybrid entanglement. For all these cases and both concrete examples of geometries, we compared the ideal scenario with situations when optical losses and spontaneous emission are taken into account.
Beyond the specific examples, our approach allows for analyzing quantum interference in other geometries, tailored to different, even continuous-in-time quantum simulation tasks. One advantage of using a cascaded configuration is that the evolution itself creates interesting kinds of entanglement from an initially non-entangled state. Thus, we presented a flexible and versatile model that can be used to characterize scalable light-matter networks for interfacing different physical platforms in quantum technologies.
@@ Acknowledgements.
CDF would like to thank the TQS group for the kind hospitality at Paderborn University. JS and LA acknowledge financial support through Ministry of Culture and Science of the State of North Rhine-Westphalia (Project PhoQC).

### v2

@@ V.1 Entanglement witnesses
One experiment-friendly way to assess entanglement is given in terms of measurable operators, dubbed entanglement witnesses [86, 87, 88, 89]. One way to construct such witnesses is based on finding the maximal (likewise, minimal) expectation values of an observable $\hat{L}$ for separable states via generalized eigenvalue problems [90, 91], the so-called separability eigenvalue equations. The thereby constructed witnesses are directly applicable to detecting multipartite entanglement in experiments [92, 93, 94].
@@
where $\hat{\rho}_{\text{sep}}$ is a nonentangled state and $\mathcal{S}$ denotes the set of all pure separable states. We have indicated with $g_{\max}$ the largest possible expected value of $\hat{L}$ for separable states. This maximal expectation value is identical to the maximal separability eigenvalue from the construction approach mentioned before [90, 91].
@@
where $\lambda_{\max}$ denotes the maximal expected value for all separable and inseparable states, i.e., the maximal ordinary eigenvalue of $\hat{L}$, and is, for all examples considered here, $\lambda_{\max}=1$. Thus, $E$ yields a value between zero and one and is zero for separable states, hence quantifying the witnessed entanglement [95].
@@
where the phases $\varphi_{k}$ are adjusted such that they coincide with the single-excitation state under study for maximizing the expectation value $\langle\hat{L}\rangle$. Regardless of the phases, the maximum expected value for separable states, i.e., maximal separability eigenvalue, reads [96, 97]
@@
With the observable described thus far, we probe how similar the system’s state is to an $N$-partite, maximally multipartite-entangled state in a $W$ configuration that includes one excitation. For the states produced in our quantum walk scenario, i.e., Eqs. (9) and (12), we can directly compute the sought-after expectation values. Since we have $N=N_{1}N_{2}$ atoms and cavities for the considered geometries, for the cavity-only entanglement, we find
@@
also yielding $\lambda_{\max}=1$. Note that $|\vec{0}\rangle$ is free of excitations and, thus, perpendicular to the single-excitation cavity state $|\vec{\alpha}\rangle$ and the single-excitation atomic state $|\vec{\beta}\rangle$. As $|\Phi\rangle$ is thus given in its Schmidt decomposition, the bipartite cavity-atom separability bound $g_{\max}$ can be readily found via the maximal Schmidt coefficient [90],
@@
Figure 4: Evolution of the entanglement in the quantum walk for a network with toroidal structure (top row) and a Möbius strip (bottom row) in terms of the entanglement quantifier in Eq. (22). Lines (in bright color) represent the ideal, lossless entanglement evolution ($\Gamma^{(C)}/\kappa=\Gamma^{(A)}/\kappa=0$) whereas decaying curves (in gray) show attenuated situations ($\Gamma^{(C)}/\kappa=\Gamma^{(A)}/\kappa=0.0025$). The multipartite entanglement of the full system is shown in (a1) and (b1) for the two geometries. The multipartite entanglement between the cavity modes alone is shown in (a2) and (b2), and the analogous case for atoms in (a3) and (b3). In (a4) and (b4), the bipartite light-matter entanglement is shown. The other parameters—in addition to the decay rates—are the same as in Fig. 3.
@@ V.2 Entanglement evolution
The evolution of the aforementioned four kinds of entanglement is displayed in Fig. 4 for the torus (top row) and Möbius strip (bottom row). In the four depicted scenarios, using $E$ as given in Eq. (22), we consider both a lossless evolution (brighter colors) and an evolution subject to attenuations (gray). The multipartite entanglement of the full system is shown in the first column. This includes all quantum correlations between the $2N$ elements (all cavity modes and atoms) and is $E=1$ for a maximal entanglement in a $W$-state configuration of the single excitation. The second column similarly depicts the entanglement between all cavity fields alone, being the $W$ entanglement of the system after tracing over the atomic degrees of freedom. Likewise, the entanglement between the atoms is shown in the third column. Finally, the bipartite light-matter entanglement, i.e., the entanglement between the collection of all atoms and the collection of all cavity fields, is presented in Figs. 4(a4) and 4(b4).
When there are no optical losses and no spontaneous decays in the system, we can see in the fourth column of Fig. 4 how the entanglement between light and matter is constant and maximum, $E(t)=1$, except for the initial time, $E(t_{0}=0)=0$, as we prepare the system in the separable configuration $|\bar{\psi}(t_{0})\rangle=|1_{k}^{(C)}\rangle$, with $k=(3,3)$. This discontinuous behavior reflects the fact that as soon as the evolution of the initial photon excitation gives a non-zero probability of finding an atomic excitation, the light and matter subsystems become entangled. When losses are included, however, these perfect correlations decay, as seen in Figs. 4(a4) and 4(b4).
The multipartite entanglement of the full system, shown in Figs. 4(a1) and 4(b1), is the most robust kind of entanglement against losses since it is the one with the highest degree of separation, i.e., each atom and cavity mode is considered as one subsystem. Note that even if concrete values are not identical for both geometries, the general behavior is analogous for the quantum walks on torus and Möbius-strip graphs, including the time decay due to losses, for all kinds of entanglement. Furthermore, we observe several instances of increasing and decreasing entanglement, including so-called sudden deaths and sudden births [98]. In general, for the four kinds of entanglement analyzed here, one can observe that a considerable amount of entanglement is created during the evolution and sustained over time, as described by the envelopes of $E(t)$ and notwithstanding short-time oscillations within.
@@ VI Conclusions
The performance of a multipartite light-matter network with arbitrary configurations as a quantum walk has been considered. The quantum dynamics and the multipartite entanglement have been analyzed, focusing on two highly nontrivial geometries: a torus as a genus-one topology and a Möbius strip as a nonorientable manifold. Clear signatures of quantumness—distinct from the dispersive random-walk evolution—were observed in our computation, which was based on the exact solutions we derived. Also, the impact of optical losses and spontaneous, atomic decays were included in our model.
Quantum interference patterns in the probability distributions were found for both geometries. This included oscillatory features in the optical part and in the atomic part, as well as between them. A comparison of random walks and quantum walks further showed that the quantum interference is highly sensitive to the underlying geometry, while no distinctive features were observable in the classical diffusion.
Additional quantum properties of the network are confirmed by the entanglement created through the open-system evolution. The analysis of entanglement includes the multipartite entanglement of the full system, the multipartite entanglement between the atoms, the multipartite entanglement between the cavity modes, and the bipartite light-matter hybrid entanglement. For all these cases and both concrete examples of geometries, we compared the ideal scenario with situations when optical losses and spontaneous emissions are taken into account.
Beyond the specific examples, our approach allows for analyzing quantum interference in other geometries, tailored to different, even continuous-in-time quantum simulation tasks. One advantage of using a cascaded configuration is that the evolution itself creates interesting kinds of entanglement from an initially nonentangled state. In addition to the system considered here, arrays of optomechanical systems present a promising platform in which our approach may be valuable for implementing quantum walks [99, 100]. Thus, we presented a flexible and versatile model that can be used to characterize scalable light-matter networks for interfacing different physical platforms in quantum technologies.
@@ Acknowledgements.
C.D.F. would like to thank the TQS group for the kind hospitality at Paderborn University. J.S. and LA acknowledge financial support through the Ministry of Culture and Science of the State of North Rhine-Westphalia (Project PhoQC).